A glass block of 10 000cm3

volume is heated from 25oC to 40oC. The volume is found to

increase by 4cm3

Calculate the coefficient of cubical expansion of glass.


1
Expert's answer
2021-10-04T10:07:07-0400

A temperature change "\Delta T" causes a change in any linear dimension "L_0" of a solid body. The change ΔL is approximately proportional to "L_0" and ΔT. Similarly, a temperature change causes a change ΔV in the volume "V_0" of any solid or liquid; ΔV is approximately proportional to "V_0" and ΔT. The quantities "\alpha" and "\beta" are the coefficients of linear expansion and volume expansion, respectively. For solids, "\beta=3\alpha=\frac{\Delta V}{V_0\,\Delta T}".


Using "V_0=10000\,{cm}^3=10^4\,{cm}^3", "\Delta T = (40-25)°C=(313.15-298.15)K=15\,K" and "\Delta V=4\,{cm^3}", and after substitution we determine the coefficient of volume expansion:


"\beta_{glass}=\frac{(4\,{ \cancel{{cm}^3}} )}{(10^4\,\cancel{{cm}^3})(15\,K)}
\\ \implies \beta_{glass}=2.667\times 10^{-5}\,K^{-1}"


In conclusion, we find the coefficient of cubical expansion of glass to have the value "\beta_{glass}=2.667\times 10^{-5}\,{K}^{-1}".

Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.